Multi-part question:

Please show explanation and step by step work!
Answer #1
Let the number of travelers be 100. Then we can represent the data in the form of the following table:
Business Personal Total
Major Airline 50% of 80 = 40 80 - 40 = 40 80% of 100 = 80
Private Owned 70% of 20 = 14 20 - 14 = 6 20% of 100 = 20
Total 54 46 100

(a) P(Business) = 54/100 = 0.54

(b) P(Business and Private Owned) = 14/100 = 0.14

(c) P(Major Airline | Business) = P(Major Airline and Business) / P(Business) = (40/100) / (54/100) = 40/54 = 0.741.